OpenStudy (anonymous):

alright, i know that the integral of 1/(1+x^2)dx is arctanx, but can someone please help me understand why?

OpenStudy (amistre64):

youd have to work it backwards from the derivative

OpenStudy (amistre64):

\[tan^{-1}(x)=y\] \[x=tan(y)\] \[1=sec^2(y)\ y'\] \[\frac{1}{sec^2{y}}= y'\]

OpenStudy (amistre64):

since tan(y) = x tan^2(y)+1 = sec^2(x) 1/x^2+1

OpenStudy (amistre64):

barring typos of course ;)

OpenStudy (anonymous):

Integral is also known as ante-derivative hence the above is true.. also integral of 1/(1+x^2)dx is *not* arctanx, you missed the constant which means a family of functions.

OpenStudy (anonymous):

right, yes my bad arctan+C. but if i saw 1/(1+x^2) on my test and had to integrate it and did not remember that it was arctan+C, there must be a way to figure it out right?

OpenStudy (amistre64):

well, since x^2 + 1 is not decomp across the reals .... maybe something to do with a complex trick

OpenStudy (amistre64):

or, maybe a usub into x = tan(u) type thing

OpenStudy (amistre64):

but then what would you do with a 1/sec^2(u) ...

OpenStudy (anonymous):

ok, i do think that i am supposed to substitute tanu=x, only because someone else told me to.

OpenStudy (amistre64):

cos^2(u) has no sin(u) to work with ....

OpenStudy (anonymous):

but then you have 1/1+tan^2u (sec^2u)

OpenStudy (amistre64):

the hard part is trying to get to tan-1 without going into complex definitions; and I dont think that is possible at the moment

OpenStudy (anonymous):

and those cancel out and just give you integral du....but i dont think that is right?

OpenStudy (amistre64):

its not right ...

OpenStudy (anonymous):

@amistre64:It's not that hard just think right angle triange :)

OpenStudy (amistre64):

triangles are pointy, and hurt lol

OpenStudy (anonymous):

haha, alright. so maybe i will just have to accept that i need to memorize it :)

OpenStudy (amistre64):

but i think the caveat is to manipulate it without remembering stuff

OpenStudy (amistre64):

its not hard to memorize :) its one of the basics for these things

OpenStudy (anonymous):

No need to memorize .. you may think of tangents of hyperbola if you want .. geometrical understanding is very essential in calculus :)

OpenStudy (amistre64):

memorize it fer sure; you might need to forget a little algebra to fit it in, but its worth it in the long run ;)

OpenStudy (amistre64):

states and their capitals are pretty useless; might try fitting it in there

OpenStudy (anonymous):

@amistre64:Memorization is bad ...... it's makes you memory dependent :P

OpenStudy (anonymous):

states capital ??????

OpenStudy (anonymous):

LOL i am to the point where i have to forget other things to fit new math in my brain. But ok, i wish i had the geometric understanding to see it...

OpenStudy (anonymous):

tangents of hyperbola? im not really sure what you mean

OpenStudy (amistre64):

sinh, cosh, and .... tanth?

OpenStudy (amistre64):

\[sinh = \frac{e^{-x}+e^x}{2}\] maybe

OpenStudy (anonymous):

x^2-y^2 = a^2 is a hyperbola ..

OpenStudy (anonymous):

Hm, alright. i know what a hyperbola is but im not sure what to do with that. It's ok tho i will continue to study. Thank you! :)